OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. : x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

@David27

OpenStudy (anonymous):

first you would place 125,000 into the y spot

OpenStudy (david27):

is given to you = 125,000 Solve for x by plugging in y: x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 x = 93,750 = Number of votes Maya received

OpenStudy (anonymous):

wow ok nvm

OpenStudy (anonymous):

lol thnx guys :)

OpenStudy (anonymous):

didnt help much

OpenStudy (david27):

your welcome

OpenStudy (anonymous):

you wer going to tho :)

OpenStudy (anonymous):

sorry i wasnt quick enough

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

david helped :)

OpenStudy (teddyiswatshecallsme):

Either i'm dumb as hell or i haven't learned this stuff. Or both. Probably both.

OpenStudy (anonymous):

hahahah its ok lol

OpenStudy (teddyiswatshecallsme):

I don't know why you're tagging me if David is so proficient in this stuff. It seems like he has this stuff on lock.